# Magnon Correlation Enables Spin Injection, Dephasing, and Transport in Canted Antiferromagnets

### v1

@@ Abstract
${\cal M}_{\alpha\beta}({\bf k},{\bf k^{\prime}})=\begin{pmatrix}{{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}^{(1)}_{{\bf q}}}\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}+\hbar\omega_{1,{\bf q}})\\ {{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}^{(2)}_{{\bf q}}}\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}+\hbar\omega_{2,{\bf q}})\\ {{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}_{-{\bf q}}^{(1)\ast}}\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}-\hbar\omega_{1,-{\bf q}})\\ {{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}_{-{\bf q}}^{(2)\ast}}\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k^{\prime}}}-\hbar\omega_{2,-{\bf q}})\end{pmatrix}$
@@
×$\times\begin{pmatrix}{\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}_{{\bf q}}^{(1)\ast},&{\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}_{{\bf q}}^{(2)\ast},&-{\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}^{(1)}_{-{\bf q}},&-{\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}^{(2)}_{-{\bf q}}\end{pmatrix}.$
@@
The electron population $f(\varepsilon_{{\bf k},\beta})=f_{0}(\varepsilon_{{\bf k}}-\varepsilon_{F})-\delta\mu_{\beta}{\partial f_{0}}/{\partial\varepsilon_{{\bf k}}}$ deviates from the equilibrium $f_{0}(\varepsilon_{{\bf k}}-\varepsilon_{F})=\{{\rm exp}[(\varepsilon_{{\bf k}}-\varepsilon_{F})/(k_{B}T)]+1\}^{-1}$ by a spin accumulation $\delta\mu_{\beta}$ that is much smaller than the Fermi energy $\varepsilon_{F}$, which exchanges the electron spin with the magnon population and correlation. Magnon population and correlation relax to their equilibrium $\rho_{\bf q}^{(0)}$ due to the magnon-phonon interaction within the time $\tau_{m}$ [16, 60] via $\partial_{t}{\rho}_{\bf q}|_{\rm rex}=-(\rho_{\bf q}-\rho_{\bf q}^{(0)})/\tau_{m}$. The steady-state solution is obtained by setting $\partial_{t}{\rho}_{\bf q}=0$.
@@
Figure 3: (a) Exchange asymmetry $\Delta$ and magnetic-field dependencies of injected magnon correlation density. (b) Distribution of the injected magnon correlation and population of mode “1” in the Brillouin zone. (c) Injected spin density $S_{z}$ along the Néel vector. (d) Calculated dephasing time $\tau_{\bf q}$ due to the magnon-electron interaction, showing excellent agreement between numerical and analytical calculations.
The spin density $S_{z}$ along the Néel vector is solely carried by magnon correlation. Indeed, as compared by Fig. 3(c) and (a), the injected spin density $S_{z}$ is proportional to $C_{12}$. The net spin injection is suppressed by decreasing the asymmetry $\Delta$ and increasing the magnetic field $H$. This can be attributed to the spin dephasing due to the transfer of magnon spin to the conductor and the intrinsic torque exerted on magnon correlation, which restricts spin injection, as analyzed below.
Focusing on spin dephasing due to magnon-electron coupling, we set $\mu_{\uparrow}=\mu_{\downarrow}$ in Eq. (7). Non-equilibrium magnon correlations now transfer spin angular momentum back to electrons. Disregarding two-magnon excitations $\langle\hat{b}^{\dagger}_{\bf q}\hat{b}^{\dagger}_{-\bf q}\rangle$, the magnon correlation $\rho_{12}({\bf q})$ obeys
@@
Gated nonlocal spin transport—The injected magnon correlations carry net spins along the Néel vector and diffuse away from the injector located around $z=0$, resulting in a spin current as depicted in Fig. 4(a). In the steady state, the diffusion equation [50]
$D_{\bf q}\partial^{2}_{z}\rho_{12}(z,{\bf q})-{\rho_{12}(z,{\bf q})}/{\tau_{\rm eff}({\bf q})}=0$ (10)
is governed by the diffusion coefficient $D_{\bf q}=v_{z}^{2}({\bf q})\tau_{\rm eff}({\bf q})$ and the effective dephasing rate $1/\tau_{\rm eff}({\bf q})=1/\tau_{m}-i\delta\omega_{\bf q}$, with $v_{z}({\bf q})=(v_{1,z}({\bf q})+v_{2,z}({\bf q}))/2$ being the group velocity of magnon correlation. With the correlation accumulation $\rho^{(0)}_{12}({\bf q})$ beneath the injector and according to Eq. (4a), Eq. (10) leads to the spin-current density in the region $z>0$ [50]
@@
Figure 4: Nonlocal spin transport tunable by the magnetic fields and metallic gate. (a) Schematic of a nonlocal spin transport device. (b) Diffusive spin current density $J_{z}(z)$ along the $\hat{\bf z}$-direction. Inset: Hanle oscillations at high magnetic field $\mu_{0}H=5$ T. (c) Gated spin transport by normal metals. (d) Spin diffusion length with and without metallic gates, also showing magnetic-field tunability.
@@
Figure 4(b) plots the exponential decay of spin-current densities $J_{z}(z)$ along the transport $\hat{\bf z}$-direction at a small magnetic field $\mu_{0}H=0.5$ T; when the magnetic field is sufficiently strong (e.g., $\mu_{0}H=5$ T), the spin current becomes oscillating as shown in the inset of Fig. 4(b) since in this case $|\delta\omega_{\bf q}|\gg 1/\tau_{m}$, such that $J_{z}(z)$ shows a pronounced Hanle oscillation. This is consistent with experimental observations [31]. As addressed by Eq. (9), the passive metallic gate is a source of spin dephasing. To detect such spin dephasing, we propose to exploit the normal metals to gate the nonlocal spin transport [19] in a device illustrated in Fig. 4(c). The random transfer of spin between normal metals and AFM brakes the spin flow by decreasing the spin diffusion length $\lambda_{s}$ at low magnetic fields through the modified effective dephasing rate $1/\tau_{\rm eff}({\bf q})\rightarrow 1/\tau_{m}+1/\tau_{\bf q}-i\delta\omega_{\bf q}$, as shown in Fig. 4(d). Nevertheless, at higher magnetic fields, the free-induction decay $\delta\omega_{\bf q}$ dominates the spin transport, so the spin flow becomes immune to the metallic gate.
@@
Discussion and conclusion.—The build-up of observable quantum coherence between magnon modes and their transport in thermodynamics is surprising in that no definite phase difference between two modes is promised. This is rooted in a mixing excitation of both modes in a quantum spin-flip process, which creates a magnon correction. The injected magnon correlation oscillates with the frequency difference of two modes, which interfere with each other for different wave vectors. The adjacent metal plays a twofold role: it injects magnon correlation but also causes dephasing of magnon correlation due to their backaction on electron spin dynamics. Our quantum theory establishes a comprehensive understanding of phase-coherent magnon-pair formation in canted AFMs, predicts a practical electrical scheme for manipulating coherent spin signals in thermal magnon transport, and applies to $\alpha$-Fe2O3 [30, 31, 32], FePS3 [68], Cr2O3 [6, 69], and many other noncollinear AFMs.

### v2

@@ III.3 Electrical spin injection and dephasing
${A}_{2\times 2}({\bf k},{\bf k^{\prime}})=\begin{pmatrix}|{{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}^{(1)}_{{\bf q}}}|^{2}\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}+\hbar\omega_{1,{\bf q}})&({{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}^{(1)}_{{\bf q}}})({\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}_{{\bf q}}^{(2)\ast})\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}+\hbar\omega_{1,{\bf q}})\\ ({{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}^{(2)}_{{\bf q}}})({\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}_{{\bf q}}^{(1)\ast})\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}+\hbar\omega_{2,{\bf q}})&|{{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}^{(2)}_{{\bf q}}}|^{2}\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}+\hbar\omega_{2,{\bf q}})\end{pmatrix}$
and
@@
${B}_{2\times 2}({\bf k},{\bf k^{\prime}})=\begin{pmatrix}-|{{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}_{-{\bf q}}^{(1)\ast}}|^{2}\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}-\hbar\omega_{1,-{\bf q}})&-({{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}_{-{\bf q}}^{(1)\ast}})({\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}^{(2)}_{-{\bf q}})\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}-\hbar\omega_{1,-{\bf q}})\\ -({{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}_{-{\bf q}}^{(2)\ast}})({\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}^{(1)}_{-{\bf q}})\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}-\hbar\omega_{2,-{\bf q}})&-|{{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}_{-{\bf q}}^{(2)\ast}}|^{2}\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}-\hbar\omega_{2,-{\bf q}})\\ \end{pmatrix}$,
the diagonal terms are related to the excitation of magnon population in modes “1” and “2”, while the off-diagonal terms drive the correlation $\langle\hat{b}_{1,{\bf q}}^{\dagger}\hat{b}_{2,{\bf q}}\rangle$ and $\langle\hat{b}_{2,{\bf q}}^{\dagger}\hat{b}_{1,{\bf q}}\rangle$ between modes “1” and “2”. In the blocks
${C}_{2\times 2}({\bf k},{\bf k^{\prime}})=\begin{pmatrix}-({{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}^{(1)}_{{\bf q}}})({\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}^{(1)}_{-{\bf q}})\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}+\hbar\omega_{1,{\bf q}})&-({{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}^{(1)}_{{\bf q}}})({\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}^{(2)}_{-{\bf q}})\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}+\hbar\omega_{1,{\bf q}})\\ -({{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}^{(2)}_{{\bf q}}})({\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}^{(1)}_{-{\bf q}})\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}+\hbar\omega_{2,{\bf q}})&-({{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}^{(2)}_{{\bf q}}})({\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}^{(2)}_{-{\bf q}})\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}+\hbar\omega_{2,{\bf q}})\end{pmatrix}$
and
${D}_{2\times 2}({\bf k},{\bf k^{\prime}})=\begin{pmatrix}({{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}_{-{\bf q}}^{(1)\ast}})({\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}_{{\bf q}}^{(1)\ast})\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}-\hbar\omega_{1,-{\bf q}})&({{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}_{-{\bf q}}^{(1)\ast}})({\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}_{{\bf q}}^{(2)\ast})\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}-\hbar\omega_{1,-{\bf q}})\\ ({{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}_{-{\bf q}}^{(2)\ast}})({\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}_{{\bf q}}^{(1)\ast})\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}-\hbar\omega_{2,-{\bf q}})&({{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}_{-{\bf q}}^{(2)\ast}})({\boldsymbol{\sigma}}_{\beta\alpha}\cdot{\boldsymbol{\chi}}_{{\bf q}}^{(2)\ast})\delta(\varepsilon_{{\bf k}}-\varepsilon_{{\bf k}^{\prime}}-\hbar\omega_{2,-{\bf q}})\end{pmatrix}$,
@@
the diagonal terms govern the generation of intraband magnon pairs $\langle\hat{b}^{\dagger}_{\xi,{\bf q}}\hat{b}^{\dagger}_{\xi,-{\bf q}}\rangle$ and $\langle\hat{b}_{\xi,{\bf q}}\hat{b}_{\xi,-{\bf q}}\rangle$, while the off-diagonal terms drive the interband pairs $\langle\hat{b}^{\dagger}_{\xi,{\bf q}}\hat{b}^{\dagger}_{\xi^{\prime},-{\bf q}}\rangle$ and $\langle\hat{b}_{\xi,{\bf q}}\hat{b}_{\xi^{\prime},-{\bf q}}\rangle$. The electron population $f(\varepsilon_{{\bf k},\beta})=f_{0}(\varepsilon_{{\bf k}}-\varepsilon_{F})-\delta\mu_{\beta}{\partial f_{0}}/{\partial\varepsilon_{{\bf k}}}$ deviates from the equilibrium $f_{0}(\varepsilon_{{\bf k}}-\varepsilon_{F})=\{{\rm exp}[(\varepsilon_{{\bf k}}-\varepsilon_{F})/(k_{B}T)]+1\}^{-1}$ by a spin accumulation $\delta\mu_{\beta}$ that is much smaller than the Fermi energy $\varepsilon_{F}$, which exchanges the electron spin with the magnon population and correlation. Magnon population and correlation relax to their equilibrium $\rho_{\bf q}^{(0)}$ due to the magnon-phonon interaction within the time $\tau_{m}$ [16, 65] via
$\partial_{t}{\rho}_{\bf q}|_{\rm rex}=-(\rho_{\bf q}-\rho_{\bf q}^{(0)})/\tau_{m}.$ (41)
The steady-state solution is obtained by setting $\partial_{t}{\rho}_{\bf q}=0$. We solve the quantum kinetic equation in the linear response regime, see Appendix B.
We have established a complete form of the quantum kinetic equation by considering the evolution of a full density matrix. Tang et al. recently considered the quantum kinetic equation for a subspace of the density matrix [63], in which the Hanle effect is the primary focus. Besides Hanle effect, our work focuses on analyzing the microscopic origins of decoherence in the spin injection and transport: 1) Intrinsic decoherence due to the free induction decay resulting from the superposition of different oscillation frequencies; 2) Extrinsic decoherence due to spin exchange with electrons in adjacent metals, based on which we propose to gate the coherent spin transport by a metallic gate.
@@
Figure 5: (a) Exchange asymmetry $\Delta$ and magnetic-field dependencies of injected magnon correlation density. (b) Distribution of the injected magnon correlation and population of mode “1” in the Brillouin zone. (c) Time evolution of magnon correlation ${\rm Re}\rho_{12,{\bf q}}$ (the blue curve) and the fitting via the exponential decay $\sim e^{-t/\tau_{\bf q}}$ (the red curve). (d) Calculated dephasing time $\tau_{\bf q}$ due to the magnon-electron interaction, showing excellent agreement between numerical and analytical calculations.
The spin density $S_{z}\approx(1/S_{0})2{\rm Re}\sum_{\bf q}\rho_{12}({\bf q})$ along the Néel vector is solely carried by magnon correlation. The net spin injection is suppressed by decreasing the asymmetry $\Delta$ and increasing the magnetic field $H$. This can be attributed to the spin dephasing due to the transfer of magnon spin to the conductor and the intrinsic torque exerted on magnon correlation, which restricts spin injection, as analyzed below.
Focusing on spin dephasing due to magnon-electron coupling, we set $\mu_{\uparrow}=\mu_{\downarrow}$ in Eq. (39). Non-equilibrium magnon correlations now transfer spin angular momentum back to electrons. The dephasing time $\tau_{\bf q}$ in the numerical result is obtained by the fitted evolution of the magnon correlation ${\rm Re}\rho_{12,{\bf q}}(t)$ by the exponential function $\sim e^{-t/\tau_{\bf q}}$, as shown in Fig. 5(c). In the numerical calculation, the overall additional relaxation $\tau_{m}$ is disregarded, such that the decay originates only from the dephasing. Analytically, disregarding two-magnon excitations $\langle\hat{b}^{\dagger}_{\bf q}\hat{b}^{\dagger}_{-\bf q}\rangle$, the magnon correlation $\rho_{12}({\bf q})$ obeys
@@
IV Gated nonlocal spin transport
The injected magnon correlations carry net spins along the Néel vector and diffuse away from the injector located around $z=0$, resulting in a spin current as depicted in Fig. 6(a). For the density matrix, the diffusion term reads
$\partial_{t}\rho_{\bf q}|_{\rm diff}=(\boldsymbol{\cal V}_{\bf q}\sigma_{3}\cdot\nabla_{\bf r}\rho_{\bf q}+\nabla_{\bf r}\rho_{\bf q}\cdot\sigma_{3}\boldsymbol{\cal V}_{\bf q})/2.$ (44)
Accordingly, in the steady state, the diffusion along the $\hat{\bf z}-$direction away from the injector is governed by
$-v_{z}({\bf q})\frac{\partial\rho_{12}(z,{\bf q})}{\partial z}-\left(\frac{1}{\tau_{m}}-i\delta\omega_{\bf q}\right)\rho_{12}(z,{\bf q})=0,$ (45)
where $v_{z}({\bf q})=(v_{1,z}({\bf q})+v_{2,z}({\bf q}))/2$ is the group velocity of the magnon correlation along the $\hat{\bf z}-$direction and $\delta\omega_{\bf q}=\omega_{1,{\bf q}}-\omega_{2,{\bf q}}$ is the frequency difference of the two magnon modes. By substituting $\rho_{12}(z,{\bf q})=-v_{z}\tau_{\rm eff}({\bf q})\partial_{z}\rho_{12}(z,{\bf q})$ back to Eq. (45) with an effective dephasing rate $1/\tau_{\rm eff}({\bf q})=1/\tau_{m}-i\delta\omega_{\bf q}$, we arrive at the second-order diffusion equation
$D_{\bf q}\partial^{2}_{z}\rho_{12}(z,{\bf q})-{\rho_{12}(z,{\bf q})}/{\tau_{\rm eff}({\bf q})}=0,$ (46)
which is governed by the diffusion coefficient $D_{\bf q}=v_{z}^{2}({\bf q})\tau_{\rm eff}({\bf q})$ and the effective dephasing rate $1/\tau_{\rm eff}({\bf q})=1/\tau_{m}-i\delta\omega_{\bf q}$, with $v_{z}({\bf q})=(v_{1,z}({\bf q})+v_{2,z}({\bf q}))/2$ being the group velocity of magnon correlation. Equation (46) has the general solution
$\rho_{12}(z>0,{\bf q})$ $=\alpha_{\bf q}\exp\left(-\frac{z}{|v_{z}({\bf q})|\tau_{\rm eff}({\bf q})}\right),$
$\rho_{12}(z<0,{\bf q})$ $=\beta_{\bf q}\exp\left(\frac{z}{|v_{z}({\bf q})|\tau_{\rm eff}({\bf q})}\right).$ (47)
At $z=0$, the continuity requires the coefficients $\alpha_{\bf q}=\beta_{\bf q}=\rho^{(0)}_{12}({\bf q})$, where $\rho^{(0)}_{12}({\bf q})$ is given by the steady-state injected magnon correlation beneath the injector.
With the correlation accumulation $\rho^{(0)}_{12}({\bf q})$ beneath the injector and according to Eq. (27a), Eq. (46) leads to the spin-current density in the region $z>0$
@@ IV Gated nonlocal spin transport
Figure 6: Nonlocal spin transport tunable by the magnetic fields and metallic gate. (a) Schematic of a nonlocal spin transport device. (b) Diffusive spin current density $J_{z}(z)$ along the $\hat{\bf z}-$direction. (c) Hanle oscillations at high magnetic field. (d) Spin current of fixed position $z$ as a function of the magnetic field as a signature of the Hanle effect. (e) Gated spin transport by normal metals. (f) Spin diffusion length with and without metallic gates, also showing magnetic-field tunability.
Figure 6(b) plots the exponential decay of spin-current densities $J_{z}(z)$ along the transport $\hat{\bf z}-$direction at a small magnetic field $\mu_{0}H=0.5$ T; when the magnetic field is sufficiently strong, the spin current becomes oscillating as shown in Fig. 6(c) since in this case $|\delta\omega_{\bf q}|\gg 1/\tau_{m}$, such that $J_{z}(z)$ shows a pronounced Hanle oscillation. This is consistent with experimental observations [31], which discusses the Hanle effect in easy-plane antiferromagnets, with a peak signal arising at a large magnetic field $\sim 8$ T. We predict the Hanle effect in canted easy-axis antiferromagnets, where the spin-current peak appears at low magnetic fields, as shown in Fig. 6(d) for fixed position. The physical origin of the Hanle effect in both descriptions is essentially the same. In the former pseudospin description [31, 33], this is interpreted as the precession of a pseudospin around a pseudo-magnetic field that is governed by the frequency difference between two modes. In our quantum description, the magnon correlation oscillates in the frequency difference between the two magnon modes. The pseudospin description employs an approximation by using an ensemble-averaged frequency difference, whereas the quantum description applies no approximation by using a full numerical calculation.
@@
As addressed by Eq. (43), the passive metallic gate is a source of spin dephasing. To detect such spin dephasing, we propose to exploit the normal metals to gate the nonlocal spin transport [19] in a device illustrated in Fig. 6(e). The random transfer of spin between normal metals and AFM brakes the spin flow by decreasing the spin diffusion length $\lambda_{s}$ at low magnetic fields through the modified effective dephasing rate $1/\tau_{\rm eff}({\bf q})\rightarrow 1/\tau_{m}+1/\tau_{\bf q}-i\delta\omega_{\bf q}$, as shown in Fig. 6(f). Nevertheless, at higher magnetic fields, the free-induction decay $\delta\omega_{\bf q}$ dominates the spin transport, so the spin flow becomes immune to the metallic gate. The efficiency of gating strongly depends on the asymmetry $\Delta$ of the coupling of the sublattice magnetic moments to electrons $\propto(1-\Delta)^{2}$. Here, $\Delta=1$ corresponds to a perfectly compensated interface, while $\Delta=0$ corresponds to an uncompensated interface. Defective interfaces enhance electron-magnon scattering, resulting in a stronger gating efficiency.
V Discussion and conclusion
@@
The build-up of observable quantum coherence between magnon modes and their transport in thermodynamics is surprising in that no definite phase difference between two modes is promised. This is rooted in a mixing excitation of both modes in a quantum spin-flip process, which creates a magnon correlation. The injected magnon correlation oscillates with the frequency difference of two modes, which interfere with each other for different wave vectors. The adjacent metal plays a twofold role: it injects magnon correlation but also causes dephasing of magnon correlation due to their backaction on electron spin dynamics. Our quantum theory establishes a comprehensive understanding of phase-coherent magnon correlation formation in canted AFMs, predicts a practical electrical scheme for manipulating coherent spin signals in thermal magnon transport, and applies to $\alpha$-Fe2O3 [30, 31, 32], FePS3 [73], Cr2O3 [6, 74], and many other noncollinear AFMs.